OpenStudy (yanasidlinskiy):

Convert 45 inches to feet and then back to inches.

OpenStudy (anonymous):

what do you mean

OpenStudy (yanasidlinskiy):

I have got: 12in/1ft X 1ft/60 So my answer is 3.75 ft. The 75 is wrong. What am I supposed to do?

OpenStudy (anonymous):

3.75 feet

OpenStudy (yanasidlinskiy):

I mean, convert to feet and then back to inches..

OpenStudy (anonymous):

thats what i got

OpenStudy (yanasidlinskiy):

Where did you get 45 from? Like, how did you solve it?

TheSmartOne (thesmartone):

12 inches = 1 feet 45 inches = x feet Make a proportion: \(\sf\Large\frac{12in}{1ft}=\frac{45 in}{x~ft}\)

TheSmartOne (thesmartone):

And just cross multiply to solve for x.

OpenStudy (yanasidlinskiy):

No....You're supposed to divide 12 inches by 60.

TheSmartOne (thesmartone):

And when you change back what you got for x feet into inches, you should get 45 inches...

OpenStudy (yanasidlinskiy):

Whoops! Here's what it is:

TheSmartOne (thesmartone):

So they want the answer in feets and inches. :P

OpenStudy (yanasidlinskiy):

Yes....

TheSmartOne (thesmartone):

45 inches Can be written as 36 inches + 9 inches = 3 feet 9 inches

OpenStudy (yanasidlinskiy):

I got, 3.75 but I just don't see how it rounds to 3.9

TheSmartOne (thesmartone):

9 inches out of 12 inches = 9/12 = .75

OpenStudy (yanasidlinskiy):

Ok, still doesn't make sense but oh well. Thanks.

OpenStudy (freckles):

\[12 \text{ inches }=1 \text{ foot } \\ \frac{1 \text{ foot}}{12 \text{ inches}}=1 \text{ and/ or } \frac{12 \text{ inches }}{1 \text{ foot }} =1 \\ \text{ you can use these two fractions to convert from inches to feet} \\ \\ \text{ and also the other way around } \\ \\ \text{ so say you want to convert } 22 \text{ inches to feet } \\ 22 \text{ inches } \cdot \frac{1 \text{ ft}}{12 {inches}}=\frac{22}{12} \text{ feet} \text{ note this can be reduced } \\ \\ \text{ we can convert } \frac{22}{12} \text{ ft} \text{ to inches as well } \\ \frac{22 }{12} \frac{ft}{1} \cdot \frac{12 \text{ inches}}{1 ft}=22 \text{ inches }\]

OpenStudy (freckles):

on and I see how you want your answer 20 let's take our 22 inches how many whole feet are in 22/12 feet?

OpenStudy (freckles):

I don't know why I typed the number 20 lol

OpenStudy (freckles):

ignore that number

OpenStudy (yanasidlinskiy):

lol.

OpenStudy (yanasidlinskiy):

We all make some kind of mistakes at times.

OpenStudy (freckles):

\[22 \text{ inches }=\frac{22}{12} \text{ feet }=\frac{12+10}{12}ft=1 \text{ ft } 10 \text{ inches }\] because 22/12 has a quotient of 1 and a remainder of 10

OpenStudy (freckles):

45/12 has a quotient of ? and a remainder of?

TheSmartOne (thesmartone):

And a note in freckles example 1 feet 10 inches = 1 feet 10/12 inches when you want in decimal :P

OpenStudy (yanasidlinskiy):

Ok, so I have something similar to this...For example if I get a question that says like.. 5.2km= ______m = _______cm

OpenStudy (yanasidlinskiy):

No, wait. Let me do this one myself and you tell me if I get it right. Ok?

OpenStudy (freckles):

1 feet and 10 inches = 1 feet and 10/12 feet @TheSmartOne

OpenStudy (yanasidlinskiy):

I'd multiply 1000 * 5.2, correct?

OpenStudy (freckles):

another example 92 inches to ? ft and ? inches \[92 \text{ inches}=\frac{92}{12} feet \] the thing you want to do is mainly write 92/12 as a mix fraction the quotient is your feets part and the remainder is your inches part

OpenStudy (yanasidlinskiy):

Quit spamming!

OpenStudy (yanasidlinskiy):

I got 5200. What's my next step?

OpenStudy (freckles):

are you working on a new problem?

OpenStudy (yanasidlinskiy):

Yes...Look above...

OpenStudy (freckles):

well that is right in meters

OpenStudy (yanasidlinskiy):

So, what would I do next? Like if I'd need it in meters and centimeters, what would I do?

OpenStudy (freckles):

well you converted it to meters it also says to convert it to centimeters

OpenStudy (freckles):

5.2km= 5200 m 5.2 km=? cm

OpenStudy (freckles):

or if you want you can convert 5200m=? cm

OpenStudy (yanasidlinskiy):

520

OpenStudy (freckles):

I think that number is a bit off :p

OpenStudy (freckles):

1 meter=100 centimeter

OpenStudy (freckles):

you have 5200 m=? cm

OpenStudy (freckles):

multiply both sides of 1 meter=100 cm by 5200

OpenStudy (yanasidlinskiy):

I'm talking about like if I needed it something like this:

OpenStudy (freckles):

that new thing you wrote doesn't make sense

OpenStudy (freckles):

the first number has no units

OpenStudy (freckles):

I already get the question though

OpenStudy (freckles):

do you know 1 meter=100 centimeter?

OpenStudy (yanasidlinskiy):

Well, the first number in units is kilometers... Yes, I do know that..

OpenStudy (freckles):

\[5200(1) \text{ meter}=5200 (100) \text{ centimeter}\] are you having trouble with the multiplying part?

OpenStudy (yanasidlinskiy):

No.

OpenStudy (freckles):

like 5200(1)=5200 and 5200(100)= <--hint just add 2 zeros at the end of 5200

OpenStudy (yanasidlinskiy):

Ya, that's what I did...How would I know which number to put in which place...Like for example, what number would go into the cm spot?

OpenStudy (freckles):

5200(100) as we talked above you just need to multiply 5200(100)

OpenStudy (yanasidlinskiy):

520,000????

OpenStudy (freckles):

\[1 \text{ meter} =100 \text{ centimeter} \\ \text{ we are going to multiply both sides by 5200 } \\ \text{ since we want to know } 5200 \text{ meters}=? \text{ centimeters} \\ \\5200(1) \text{ meter}=5200 (100) \text{ centimeter}\] \[5200 \text{ meter}=520000 \text{ centimeter}\] yes

OpenStudy (yanasidlinskiy):

Oh, ok! Lol, sorry, I was confusing myself, haha! My bad. But thanks for helping me though. I get it.

OpenStudy (yanasidlinskiy):

@kanard then delete your replies. Stop yelling.